# Finite temperature detection of quantum critical points via internal quantum teleportation

### v2

@@ Abstract
We show that the teleportation protocol can be efficiently used to detect quantum critical points using finite temperature data even if all resources needed to its implementation lie within the system under investigation. Contrary to a previous proposal, there is no need to use an external qubit as the input state to be teleported to one of the qubits within the system. Here, we use a pair of nearest neighbor spins from an infinite spin-1/2 chain in equilibrium with a heat bath as the entangled resource of the quantum teleportation protocol and a third adjacent qubit within the chain itself as the input state to be teleported. For several spin chain models subjected to an external magnetic field, we show that the efficiency of the teleportation protocol is severely affected as we cross the quantum critical points associated with those spin chains. This abrupt change in efficiency gives us a clear indication of a quantum phase transition.
@@ I Introduction
In this work we show, under very general conditions, that we can simplify the tools to detect QCPs at finite $T$ given in Ref. pav23 , while still keeping their efficiency, robustness, and scalability. In the present proposal, we do not need an external qubit from the system to implement the teleportation based QCP detector. We even do not need to repeat the procedure using several different external qubits covering the whole Bloch sphere in order to detect a QCP pav23 . In the present proposal, the input state to be teleported is now fixed and belongs to the many-body system itself (see Fig. 1). Also, and similarly to the tool developed in Ref. pav23 , we do not need to know the order parameter associated with the QPT to implement it (see Refs. wer10b ; pav23 for more details).
Being more specific, we study the ability of the present tools, developed in Sec. II, to detect QCPs at finite or zero temperature for several spin-1/2 chains in the thermodynamic limit (infinite chains). We study the XXZ model subjected to a longitudinal external magnetic field in Sec. III and the Ising model and the XY model in a transverse magnetic field in Sec. IV. We end this work giving in Sec. V an extensive analysis of the theoretical and experimental resources needed to apply these tools and in Sec. VI we provide our concluding remarks.
@@ II.1 General settings
where $\Delta r_{\alpha}(t)=Tr(\rho_{1}\sigma^{\alpha})-Tr(\rho_{{}_{B_{j}}}\sigma^{\alpha})$, with $\sigma^{\alpha}$ being a Pauli matrix. Geometrically, the trace distance is half the Euclidean distance between the points on the Bloch sphere associated with the density matrices $\rho_{1}$ and $\rho_{{}_{B_{j}}}$ nie00 . Therefore, when the two states are equal we have $D_{j}=0$ and the more far apart the two states are in the Bloch sphere the greater $D_{j}$. For orthogonal pure states we have $D_{j}=1$, the maximum possible value for the trace distance.
@@
and the second QCP point detector is built minimizing $\overline{D}(S_{k})$ over all sets $S_{k}$,
@@ II.2 Specific settings
The density matrix describing three nearest neighbors from the spin chain is given by $\rho_{123}$, where it is obtained tracing out from $\varrho$ all but spins $1,2$ and $3$. Note that the initial state $\rho_{123}$ can be obtained using any three nearest neighbors from the spin chain. After putting the whole chain in contact with a heat reservoir at temperature $T$ and waiting for the system to reach the thermodynamic equilibrium, any three consecutive spins of the chain can be used to implement the internal teleportation protocol as described below.
@@
Writing $\rho$ as given by Eq. (31) readily suggests the steps Alice needs to do to effectively obtain state (1). She projects qubit 1 in the computational basis. She either obtains $|0\rangle_{1}$ or $|1\rangle_{1}$ with probabilities $p_{0}$ and $p_{1}$, respectively. According to the measurement postulate of quantum mechanics nie00 , the state $\rho_{123}$ changes in each case to
@@
Note that due to the linearity of all the steps related to the implementation of the teleportation protocol, from the experimental point of view Alice’s job to effectively obtain Eq. (1) is very simple. She just needs to measure the qubit $1$ in the computational basis and then either do nothing or implement the bit flip operation before initiating the standard teleportation protocol as described in Sec. II.1. Being more specific, if she measures $|0\rangle$ she does nothing to her qubit $(p_{0})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before starting the teleportation protocol. If she measures $|1\rangle$, she does nothing $(p_{1})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before implementing the teleportation protocol. Also, she must work with an equal amount of cases where she measured $|0\rangle$ or $|1\rangle$ to mimic the state given by Eq. (31) following the prescription given above. This is easily accomplished discarding the exceeding cases.
The bottom line is that by proceeding as described above and due to the linearity of all the steps involved in the execution of the teleportation protocol, the effective initial state will be given by Eq. (1) and Bob’s qubit at the end of several runs of the teleportation protocol will effectively be described by Eq. (2) whenever Alice measures the Bell state $j$. Figure 2 illustrates all steps of a single run of the protocol.
@@
Alternatively one could use the state $\rho_{123}$ to theoretically simulate the execution of the internal teleportation protocol. The state $\rho_{123}$ is a $8\times 8$ matrix depending not only on one- and two-point correlation functions but also on the genuine three-point ones ($\langle\sigma^{\alpha}_{j}\sigma^{\beta}_{j+1}\sigma^{\gamma}_{j+1}\rangle$). Nevertheless, due to the factorization property of the correlation functions of the XXZ spin chain boo06 , the three-point correlations for this model can be expressed very non-trivially in terms of the one- and two-point correlations. However, the computation of the three-point correlation functions for finite temperature and for the other models here investigated is a quite involved task boo06 ; boo08 ; tri10 . Thus, in order to test the usefulness of the internal teleportation QCP detector, we worked with an initial state that depends at most on two-point correlation functions of two consecutive spin. It would be very interesting to extend the present analysis to the case where the state $\rho_{123}$ is used as the initial state. We believe that the internal teleportation QCP detector will also work in this case. Perhaps giving results even better than the ones reported here. We should also note that from the experimental point of view working with $\rho_{123}$ is even simpler. In this case Alice will neither need to project qubit $1$ onto the computational basis or flip her spin before implementing the standard teleportation protocol.
@@
Remark 3: Equations (42) and (45) do not depend on the two-point correlation functions $xx=\left\langle\sigma_{j}^{x}\sigma_{j+1}^{x}\right\rangle$ and $yy=\left\langle\sigma_{j}^{y}\sigma_{j+1}^{y}\right\rangle$, as can be seen looking at Eqs. (38), (39), (43), and (44). The underlying reason for this is the specific form of $\rho_{1}$ [Eq. (28)]. Whenever the input state to be teleported is diagonal in the standard basis $\{|0\rangle,|1\rangle\}$, the expressions for the fidelity and for the trace distance will not depend on $xx$ and $yy$ if $\rho_{23}$ is given by Eq. (20). For input states having non-diagonal terms, though, a direct calculation shows that both the fidelity and the trace distance will also depend on $xx$ and $yy$.
Remark 4: Contrary to the scenario in which the qubit to be teleported is external to the spin chain pav23 , the expressions for the fidelity and for the trace distance now depend on the one-point correlation function $z=\left\langle\sigma_{j}^{z}\right\rangle$. This happens because in the present case (internal teleportation), the input state to be teleported is a function of $z$, as can be seen looking at Eq. (28). And this implies that the teleported states with Bob are, according to Eq. (2),
@@ III The XXZ model in an external field
This problem is solved working with the trace distance. In Fig. 4 we show $\overline{\mathcal{D}}$ as a function of $\Delta$ for $h=12$ and the same values of $kT$ given in Fig. 3.
@@
Looking at Fig. 6, we note that for all temperatures the extremum of the first order derivative of $\overline{\mathcal{D}}$ gives the best approximation for the exact value of the first QCP (upper panel of Fig. 6). The second QCP is also better approximated by the the extremum of the second order derivative of $\overline{\mathcal{D}}$ (lower panel of Fig. 6). However, as we decrease $kT$ the other quantities shown in Fig. 6 catch up. Around $kT=0.1$ they are all equally efficient in estimating the location of the second QCP. Also, remembering that for the external teleportation strategy the behavior of the derivatives of $\overline{\mathcal{F}}$ are equal to the behavior of the derivatives of $\langle\sigma^{z}_{j}\sigma^{z}_{j+1}\rangle$ pav23 , we have that the internal teleportation strategy gives better estimates for the location of the two QCPs when compared with the external teleportation approach.
Excluding the $kT=0$ point, we can implement linear and quadratic regressions to find the best curves fitting the remaining data ($kT=0.1,0.2,0.3,0.4,0.5$) pav23 . For all quantities given in the lower panel of Fig. 6, a linear regression is enough to predict the correct location of the QCP $\Delta_{2}$ as we take the $T=0$ limit. The predicted location of $\Delta_{2}$ is given within an accuracy of ±$\pm 0.02$, the estimated numerical error for the location of the extrema of the second order derivatives of all the quantities given in Fig. 6. The same conclusion applies for all but one of the curves given in the upper panel of Fig. 6, where again the fitted straight lines give the correct location of the QCP $\Delta_{1}$ within an accuracy of ±$\pm 0.01$, the numerical error for the locations of the extrema of the first order derivatives of all the quantities given in Fig. 6. To achieve this same level of accuracy for $\overline{\mathcal{D}}$, though, we need a quadratic regression.
@@
Remark 7: Instead of using $\overline{\mathcal{D}}$ [Eq. (45)] and its derivatives to estimate the QCPs using finite $T$ data, we could have worked directly with either $\overline{D}(S_{\Phi^{\pm}})$ or $\overline{D}(S_{\Psi^{\pm}})$ [Eqs. (43) or (44)] and their derivatives. All three quantities give similar results in predicting the location of the QCPs.
@@ IV.1 The $\lambda$ transition
Figure 8: (color online) Same as Fig. 7 but now $\gamma=0.5$.
Figure 9: (color online) Same as Fig. 7 but now $\gamma=1.0$ (Ising model in a transverse field). The upper inset highlights the $T=0$ curve of $\overline{\mathcal{F}}$ and the lower inset its derivative with respect to $\lambda$.
@@
Figure 11: (color online) Same as Fig. 10 but now $\gamma=0.5$.
Figure 12: (color online) Same as Fig. 10 but now $\gamma=1.0$ (Ising model in a transverse field). The lower inset highlights the $T=0$ curve of $\overline{\mathcal{D}}$ and the upper inset its derivative with respect to $\lambda$.
@@
If we now fix $\gamma=0.5$, the extrema of $d^{2}\overline{\mathcal{F}}/d\lambda^{2}$ lie closer to the true value of the QCP than those of $d^{2}\overline{\mathcal{D}}/d\lambda^{2}$ when $kT>0.02$ (see Fig. 14). For high values of $kT$, the extrema of $d(xx)/d\lambda$ is the optimal choice to estimate the QCP and, as we keep decreasing $T$, all quantities eventually give the exact location for the QCP. It is worth noting that if we take into account the errors (yellow-shaded regions in Fig. 14) in numerically locating the extrema of $d^{2}\overline{\mathcal{F}}/d\lambda^{2}$, for all temperatures shown in Fig. 14 the exact value of the QCP lie within the range of the possible values for the location of those extrema.
@@ V Discussion
The present proposal also retains the remarkable features of the former one pav23 . From a theoretical point of view, all we need to calculate the efficiency of the teleportation protocol, quantified by either the maximum mean fidelity $\overline{\mathcal{F}}$ or the minimum mean trace distance $\overline{\mathcal{D}}$, is the state describing two nearest neighbor qubits of the spin chain [Eq. (20)], with the latter assumed to be in equilibrium with a heat bath at temperature $T$. This two-qubit density matrix is theoretically computed tracing out all but these two nearest neighbors from the canonical ensemble density matrix describing the complete chain. Equivalently, if we are able to compute all the one- and two-point correlation functions we can determine the two-qubit density matrix describing a pair of spins since it is completely characterized by these correlation functions.
@@
As already pointed out in Ref. pav23 , some of the standard quantities used to detect QCPs (magnetization or magnetic susceptibility, for instance) become less effective to correctly identify the QCPs with finite $T$ data wer10b . And some quantum information theory quantities, in particular the entanglement of formation woo98 , are zero at and around the QCPs as we increase $T$. These quantities cannot be employed to spotlight QCPs after a certain temperature threshold wer10b . On the other hand, a specific measure of non-classical correlations between two systems, namely, quantum discord oll01 ; hen01 , is extremely robust to spotlight QCPs at finite $T$ wer10b . The present internal teleportation based tools to detect QCPs as well as the ones of Ref. pav23 share the quantum discord robustness to detect QCPs with finite $T$ data. [^2] Note that it is important to analyze the $T=0$ limit from the finite $T$ data to make sure that no spurious classical critical point is “contaminating” the finite $T$ data. Only if the finite $T$ data consistently tend to a specific critical point as we decrease $T$ we can be reassured that we are properly seeing a QCP. See, for instance, Figs. 6, 13, 14, and 15. In addition to that, and contrary to quantum discord, the present tools and the ones of Ref. pav23 do not become computationally intractable if we deal with high spin systems (high dimensional Hilbert spaces) pav23 .
@@
Turning our attention to the experimental aspects of the present proposal, it is clearly simpler than the external teleportation QCP detector tools pav23 . There is no need to bring a qubit outside from the spin chain to implement the present proposal. The input state to be teleported belongs to the chain itself. Also, the present scheme and that of Ref. pav23 have a straightforward experimental interpretation since they are built on the quantum teleportation protocol. Quantum discord, though, has no direct experimental procedure to its determination.
@@
Furthermore, to execute the internal teleportation protocol, and thus the present proposal, we need to be able to implement single qubit measurements, Bell state measurements, and local unitary operations on the qubits that belong to a spin chain. The single qubit measurement we need is a projective measurement onto the eigenstates of the $\sigma^{z}$ observable (our computational basis). A Bell state measurement is a joint measurement involving two qubits that aims to project the two qubits onto one of the four Bell states. This measurement can be divided into the following sequence of three steps nie00 . First, we implement a controlled-not (CNOT) gate on the pair of spins we will be measuring, now called control and target qubits. The CNOT is a two-qubit (non-local) gate whose “truth table” in the standard basis is such that it flips the spin of the target qubit if, and only if, the control qubit is given by $|1\rangle$. The standard basis can be, for instance, the eigenstates of the spin operator $\sigma^{z}$. Second, we implement a Hadamard gate on the control qubit of the CNOT gate. This is a single (local) qubit gate whose truth table is $(|0\rangle+(-)|1\rangle)/\sqrt{2}\rightarrow|0\rangle(|1\rangle)$. Third, we measure the two qubits in the standard (computational) basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Each one of the four previous outcomes in the standard basis implies that we projected the two qubits onto one of the four Bell states nie00 . The single qubit and Bell measurements are Alice’s job while Bob’s job is to measure his qubit after implementing the appropriate local operation (local gate) as given by Eqs. (8)-(11). All these gates, the ones needed by Alice and Bob, are already being implemented in spin chain-like systems using state of the art techniques ron15 ; bra19 ; xie19 ; noi22 ; xue22 ; mad22 ; xie22 . We believe that in the near future they can be implemented in the sequence described above to execute the teleportation protocol and consequently the present proposal.
@@
For instance, using superconducting qubits to realize a spin chain, it is already possible to experimentally prepare tens of spins simultaneously with coherence times around $100\mu s$. Also, single and two-qubit gates in this platform can be implemented within $10ns$ to $100ns$, leading to the execution of $10^{3}$ to $10^{4}$ gates per coherence time gam17 . We can also implement spin chains experimentally using quantum dots and wells. For quantum wells built with GaAs, at room temperature the relaxation time is of the order of a few nanoseconds ohn99 . And for GaAs quantum dots, the relaxation time for a qubit is about 50 $\mu s$ at $T\approx 20mK$ han03 while for arrays of Ge/Si quantum dots we have a relaxation time of about $10\mu s$ at $T\approx 5K$ zin10 . These facts, together with the fact that in silicon quantum dots at $T\approx 1K$ we can implement single and two-qubit gates in less than $100ns$ pet22 , imply that at low temperatures ($\approx 1K$) and for silicon quantum dots it is possible to execute one hundred or so gates before the spin chain thermalizes. This number of gates is at least one order of magnitude more than what we need to implement the present proposal, where only a few gates (less than ten) is needed at a given run of the teleportation protocol. [^3] The quantum circuit representing all the steps of the teleportation protocol can be seen in Ref. nie00 while a review on thermalization is given in Ref. mor18 . See also Ref. pek23 for a quantum circuit realization of a heat bath in solid quantum systems.

### v3

@@ Abstract
We show that the teleportation protocol can be efficiently used to detect quantum critical points using finite temperature data even if all resources needed for its implementation lie within the system under investigation. Contrary to a previous proposal, there is no need to use an external qubit as the input state to be teleported to one of the qubits within the system. Here, we use a pair of nearest neighbor spins from an infinite spin-1/2 chain in equilibrium with a heat bath as the entangled resource of the quantum teleportation protocol and a third adjacent qubit within the chain itself as the input state to be teleported. For several spin chain models subjected to an external magnetic field, we show that the efficiency of the teleportation protocol is severely affected as we cross the quantum critical points associated with those spin chains. This abrupt change in efficiency gives us a clear indication of a quantum phase transition.
@@ I Introduction
In this work we show, under very general conditions, that we can simplify the tools to detect QCPs at finite $T$ given in Ref. pav23 , while still keeping their efficiency, robustness, and scalability. In the present proposal, we do not need an external qubit from the system to implement the teleportation based QCP detector. We do not even need to repeat the procedure using several different external qubits covering the whole Bloch sphere in order to detect a QCP pav23 . In the present proposal, the input state to be teleported is now fixed and belongs to the many-body system itself (see Fig. 1). Also, and similarly to the tool developed in Ref. pav23 , we do not need to know the order parameter associated with the QPT to implement it (see Refs. wer10b ; pav23 for more details).
Being more specific, we study the ability of the present tools, developed in Sec. II, to detect QCPs at finite or zero temperature for several spin-1/2 chains in the thermodynamic limit (infinite chains). We study the XXZ model subjected to a longitudinal external magnetic field in Sec. III and the Ising model and the XY model in a transverse magnetic field in Sec. IV. We end this work by giving in Sec. V an extensive analysis of the theoretical and experimental resources needed to apply these tools and in Sec. VI we provide our concluding remarks.
@@ II.1 General settings
where $\Delta r_{\alpha}(t)=Tr(\rho_{1}\sigma^{\alpha})-Tr(\rho_{{}_{B_{j}}}\sigma^{\alpha})$, with $\sigma^{\alpha}$ being a Pauli matrix. Geometrically, the trace distance is half the Euclidean distance between the points on the Bloch sphere associated with the density matrices $\rho_{1}$ and $\rho_{{}_{B_{j}}}$ nie00 . Therefore, when the two states are equal we have $D_{j}=0$ and the farther apart the two states are in the Bloch sphere the greater $D_{j}$ is. For orthogonal pure states we have $D_{j}=1$, the maximum possible value for the trace distance.
@@
and the second QCP point detector is built by minimizing $\overline{D}(S_{k})$ over all sets $S_{k}$,
@@ II.2 Specific settings
The density matrix describing three nearest neighbors from the spin chain is given by $\rho_{123}$, which is obtained by tracing out from $\varrho$ all but spins $1,2$ and $3$. Note that the initial state $\rho_{123}$ can be obtained using any three nearest neighbors from the spin chain. After putting the whole chain in contact with a heat reservoir at temperature $T$ and waiting for the system to reach the thermodynamic equilibrium, any three consecutive spins of the chain can be used to implement the internal teleportation protocol as described below.
@@
Writing $\rho$ as given by Eq. (31) readily suggests the steps Alice needs to do to effectively obtain state (1). She projects qubit 1 in the computational basis. She obtains either $|0\rangle_{1}$ or $|1\rangle_{1}$ with probabilities $p_{0}$ and $p_{1}$, respectively. According to the measurement postulate of quantum mechanics nie00 , the state $\rho_{123}$ changes in each case to
@@
Note that due to the linearity of all the steps related to the implementation of the teleportation protocol, from the experimental point of view Alice’s job to effectively obtain Eq. (1) is very simple. She just needs to measure the qubit $1$ in the computational basis and then either do nothing or implement the bit flip operation before initiating the standard teleportation protocol as described in Sec. II.1. Being more specific, if she measures $|0\rangle$ she does nothing to her qubit $(p_{0})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before starting the teleportation protocol. If she measures $|1\rangle$, she does nothing $(p_{1})^{2}$ of the times or she flips it $p_{0}p_{1}$ of the times before implementing the teleportation protocol. Also, she must work with an equal number of cases where she measured $|0\rangle$ or $|1\rangle$ to mimic the state given by Eq. (31) following the prescription given above. This is easily accomplished discarding the exceeding cases.
The bottom line is that by proceeding as described above and due to the linearity of all the steps involved in the execution of the teleportation protocol, the effective initial state will be given by Eq. (1) and Bob’s qubit at the end of several runs of the teleportation protocol will effectively be described by Eq. (2) whenever Alice measures the Bell state $j$. Figure 2 illustrates all the steps of a single run of the protocol.
@@
Alternatively one could use the state $\rho_{123}$ to theoretically simulate the execution of the internal teleportation protocol. The state $\rho_{123}$ is a $8\times 8$ matrix depending not only on one- and two-point correlation functions but also on the genuine three-point ones ($\langle\sigma^{\alpha}_{j}\sigma^{\beta}_{j+1}\sigma^{\gamma}_{j+1}\rangle$). Nevertheless, due to the factorization property of the correlation functions of the XXZ spin chain boo06 , the three-point correlations for this model can be expressed very non-trivially in terms of the one- and two-point correlations. However, the computation of the three-point correlation functions for finite temperature and for the other models here investigated is a quite involved task boo06 ; boo08 ; tri10 . Thus, in order to test the usefulness of the internal teleportation QCP detector, we worked with an initial state that depends at most on two-point correlation functions of two consecutive spin. It would be very interesting to extend the present analysis to the case where the state $\rho_{123}$ is used as the initial state. We believe that the internal teleportation QCP detector will also work in this case. Perhaps giving results even better than the ones reported here. We should also note that from the experimental point of view working with $\rho_{123}$ is even simpler. In this case Alice will not need to project qubit $1$ onto the computational basis or flip her spin before implementing the standard teleportation protocol.
@@
Remark 3: Equations (42) and (45) do not depend on the two-point correlation functions $xx=\left\langle\sigma_{j}^{x}\sigma_{j+1}^{x}\right\rangle$ and $yy=\left\langle\sigma_{j}^{y}\sigma_{j+1}^{y}\right\rangle$, as can be seen by looking at Eqs. (38), (39), (43), and (44). The underlying reason for this is the specific form of $\rho_{1}$ [Eq. (28)]. Whenever the input state to be teleported is diagonal in the standard basis $\{|0\rangle,|1\rangle\}$, the expressions for the fidelity and for the trace distance will not depend on $xx$ and $yy$ if $\rho_{23}$ is given by Eq. (20). For input states having non-diagonal terms, though, a direct calculation shows that both the fidelity and the trace distance will also depend on $xx$ and $yy$.
Remark 4: Contrary to the scenario in which the qubit to be teleported is external to the spin chain pav23 , the expressions for the fidelity and for the trace distance now depend on the one-point correlation function $z=\left\langle\sigma_{j}^{z}\right\rangle$. This happens because in the present case (internal teleportation), the input state to be teleported is a function of $z$, as can be seen by looking at Eq. (28). And this implies that the teleported states with Bob are, according to Eq. (2),
@@ III The XXZ model in an external field
This problem is solved by working with the trace distance. In Fig. 4 we show $\overline{\mathcal{D}}$ as a function of $\Delta$ for $h=12$ and the same values of $kT$ given in Fig. 3.
@@
Looking at Fig. 6, we note that for all temperatures the extremum of the first order derivative of $\overline{\mathcal{D}}$ gives the best approximation for the exact value of the first QCP (upper panel of Fig. 6). The second QCP is also better approximated by the the extremum of the second order derivative of $\overline{\mathcal{D}}$ (lower panel of Fig. 6). However, as we decrease $kT$ the other quantities shown in Fig. 6 catch up. Around $kT=0.1$ they are all equally efficient in estimating the location of the second QCP. Also, remembering that for the external teleportation strategy the behavior of the derivatives of $\overline{\mathcal{F}}$ is equal to the behavior of the derivatives of $\langle\sigma^{z}_{j}\sigma^{z}_{j+1}\rangle$ pav23 , we have that the internal teleportation strategy gives better estimates for the location of the two QCPs when compared with the external teleportation approach.
Excluding the $kT=0$ point, we can implement linear and quadratic regressions to find the best curves fitting the remaining data ($kT=0.1,0.2,0.3,0.4,0.5$) pav23 . For all quantities given in the lower panel of Fig. 6, a linear regression is enough to predict the correct location of the QCP $\Delta_{2}$ as we take the $T=0$ limit. The predicted location of $\Delta_{2}$ is given within an accuracy of ±$\pm 0.02$, the estimated numerical error for the locations of the extrema of the second order derivatives of all the quantities given in Fig. 6. The same conclusion applies for all but one of the curves given in the upper panel of Fig. 6, where again the fitted straight lines give the correct location of the QCP $\Delta_{1}$ within an accuracy of ±$\pm 0.01$, the numerical error for the locations of the extrema of the first order derivatives of all the quantities given in Fig. 6. To achieve this same level of accuracy for $\overline{\mathcal{D}}$, though, we need a quadratic regression.
@@
Remark 7: Instead of using $\overline{\mathcal{D}}$ [Eq. (45)] and its derivatives to estimate the QCPs using finite $T$ data, we could have worked directly with either $\overline{D}(S_{\Phi^{\pm}})$ or $\overline{D}(S_{\Psi^{\pm}})$ [Eq. (43) or (44)] and their derivatives. All three quantities give similar results in predicting the location of the QCPs.
@@ IV.1 The $\lambda$ transition
Figure 8: (color online) Same as Fig. 7 but now $\gamma=0.5$. The dotted lines mark the QCP $\lambda_{c}$ and for the solid curves the temperature increases from top to bottom when $\lambda<\lambda_{c}$.
Figure 9: (color online) Same as Fig. 7 but now $\gamma=1.0$ (Ising model in a transverse field). The upper inset highlights the $T=0$ curve of $\overline{\mathcal{F}}$ and the lower inset its derivative with respect to $\lambda$. The dotted lines mark the QCP $\lambda_{c}$ and for the solid curves the temperature increases from top to bottom when $\lambda<\lambda_{c}$.
@@
Figure 11: (color online) Same as Fig. 10 but now $\gamma=0.5$. The inset shows the $T=0$ curve and the dotted lines mark the QCP $\lambda_{c}$. For the solid curves the temperature decreases from top to bottom when $\lambda<\lambda_{c}$.
Figure 12: (color online) Same as Fig. 10 but now $\gamma=1.0$ (Ising model in a transverse field). The lower inset highlights the $T=0$ curve of $\overline{\mathcal{D}}$ and the upper inset its derivative with respect to $\lambda$. The dotted lines mark the QCP $\lambda_{c}$. For the solid curves the temperature decreases from top to bottom when $\lambda<\lambda_{c}$.
@@
If we now fix $\gamma=0.5$, the extrema of $d^{2}\overline{\mathcal{F}}/d\lambda^{2}$ lie closer to the true value of the QCP than those of $d^{2}\overline{\mathcal{D}}/d\lambda^{2}$ when $kT>0.02$ (see Fig. 14). For high values of $kT$, the extrema of $d(xx)/d\lambda$ are the optimal choice to estimate the QCP and, as we keep decreasing $T$, all quantities eventually give the exact location for the QCP. It is worth noting that if we take into account the errors (yellow-shaded regions in Fig. 14) in numerically locating the extrema of $d^{2}\overline{\mathcal{F}}/d\lambda^{2}$, for all temperatures shown in Fig. 14 the exact value of the QCP lies within the range of possible values for the location of those extrema.
@@ V Discussion
The present proposal also retains the remarkable features of the former one pav23 . From a theoretical point of view, all we need to calculate the efficiency of the teleportation protocol, quantified by either the maximum mean fidelity $\overline{\mathcal{F}}$ or the minimum mean trace distance $\overline{\mathcal{D}}$, is the state describing two nearest neighbor qubits of the spin chain [Eq. (20)], with the latter assumed to be in equilibrium with a heat bath at temperature $T$. This two-qubit density matrix is theoretically computed by tracing out all but these two nearest neighbors from the canonical ensemble density matrix describing the complete chain. Equivalently, if we are able to compute all the one- and two-point correlation functions we can determine the two-qubit density matrix describing a pair of spins since it is completely characterized by these correlation functions.
@@
As already pointed out in Ref. pav23 , some of the standard quantities used to detect QCPs (magnetization or magnetic susceptibility, for instance) become less effective to correctly identify the QCPs with finite $T$ data wer10b . And some quantum information theory quantities, in particular the entanglement of formation woo98 , are zero at and around the QCPs as we increase $T$. These quantities cannot be employed to spotlight QCPs after a certain temperature threshold wer10b . On the other hand, a specific measure of non-classical correlations between two systems, namely, quantum discord oll01 ; hen01 , is extremely robust to spotlight QCPs at finite $T$ wer10b . The present internal teleportation based tools to detect QCPs as well as the ones of Ref. pav23 share the quantum discord robustness to detect QCPs with finite $T$ data. [^2] Note that it is important to analyze the $T=0$ limit from the finite $T$ data to make sure that no spurious classical critical point is “contaminating” the finite $T$ data. Only if the finite $T$ data consistently tend to a specific critical point as we decrease $T$ can we be reassured that we are properly seeing a QCP. See, for instance, Figs. 6, 13, 14, and 15. In addition to that, and contrary to quantum discord, the present tools and the ones of Ref. pav23 do not become computationally intractable if we deal with high spin systems (high dimensional Hilbert spaces) pav23 .
@@
Turning our attention to the experimental aspects of the present proposal, it is clearly simpler than the external teleportation QCP detector tools pav23 . There is no need to bring a qubit outside from the spin chain to implement the present proposal. The input state to be teleported belongs to the chain itself. Indeed, contrary to the external teleportaton QCP detector, there is no need to repeat the protocol with many different input qubits that cover the whole Bloch sphere pav23 . For the internal teleportation QCP detector, the input state is fixed and belongs to the system under study. Also, the present scheme and that of Ref. pav23 have a straightforward experimental interpretation since they are built on the quantum teleportation protocol. Quantum discord, though, has no direct experimental procedure to its determination.
@@
To execute the internal teleportation protocol, and thus the present proposal, we need to be able to implement single qubit measurements, Bell state measurements, and local unitary operations on the qubits that belong to a spin chain. The single qubit measurement we need is a projective measurement onto the eigenstates of the $\sigma^{z}$ observable (our computational basis). A Bell state measurement is a joint measurement involving two qubits that aims to project the two qubits onto one of the four Bell states. This measurement can be divided into the following sequence of three steps nie00 . First, we implement a controlled-NOT (CNOT) gate on the pair of spins we will be measuring, now called control and target qubits. The CNOT is a two-qubit (non-local) gate whose “truth table” in the standard basis is such that it flips the spin of the target qubit if, and only if, the control qubit is given by $|1\rangle$. The standard basis can be, for instance, the eigenstates of the spin operator $\sigma^{z}$. Second, we implement a Hadamard gate on the control qubit of the CNOT gate. This is a single (local) qubit gate whose truth table is $(|0\rangle+(-)|1\rangle)/\sqrt{2}\rightarrow|0\rangle(|1\rangle)$. Third, we measure the two qubits in the standard (computational) basis $\{|00\rangle,|01\rangle,|10\rangle,|11\rangle\}$. Each one of the four previous outcomes in the standard basis implies that we projected the two qubits onto one of the four Bell states nie00 . The single qubit and Bell measurements are Alice’s job while Bob’s job is to measure his qubit after implementing the appropriate local operation (local gate) as given by Eqs. (8)-(11). All these gates, the ones needed by Alice and Bob, are already being implemented in spin chain-like systems using state of the art techniques ron15 ; bra19 ; xie19 ; noi22 ; xue22 ; mad22 ; xie22 . We believe that in the near future they can be implemented in the sequence described above to execute the teleportation protocol and consequently the present proposal. We should remark that the experiments reported in Refs. ron15 ; bra19 ; xie19 ; noi22 ; xue22 ; mad22 ; xie22 were done using small systems. However, some of them implemented efficiently the one- and two-qubit gates at room temperature.
@@
Furthermore, using superconducting qubits to realize a spin chain, it is already possible to experimentally prepare tens of spins simultaneously with coherence times around $100\mu s$. Also, single and two-qubit gates in this platform can be implemented within $10ns$ to $100ns$, leading to the execution of $10^{3}$ to $10^{4}$ gates per coherence time gam17 . We can also implement spin chains experimentally using quantum dots and wells. For quantum wells built with GaAs, at room temperature the relaxation time is of the order of a few nanoseconds ohn99 . And for GaAs quantum dots, the relaxation time for a qubit is about 50 $\mu s$ at $T\approx 20mK$ han03 while for arrays of Ge/Si quantum dots we have a relaxation time of about $10\mu s$ at $T\approx 5K$ zin10 . These facts, together with the fact that in silicon quantum dots at $T\approx 1K$ we can implement single and two-qubit gates in less than $100ns$ pet22 , imply that at low temperatures ($\approx 1K$) and for silicon quantum dots it is possible to execute one hundred or so gates before the spin chain thermalizes. This number of gates is at least one order of magnitude more than what we need to implement the present proposal, where only a few gates (fewer than ten) are needed at a given run of the teleportation protocol. [^3] The quantum circuit representing all the steps of the teleportation protocol can be seen in Ref. nie00 while a review on thermalization is given in Ref. mor18 . See also Ref. pek23 for a quantum circuit realization of a heat bath in solid quantum systems. Furthermore, with the advent of a quantum computer, it will be possible to efficiently simulate and even prepare thermal states bra20 ; sag21 ; bra23 ; wan23 , which in turn can be used to implement the present QCP detection strategy.
Note that the time span to execute all the steps of the teleportation protocol plus Bob’s measurement of the teleported qubit should not exceed the “relaxation time” of the spin chain. That is, all the steps should be performed before the spin chain returns to its equilibrium state with the heat bath. This equilibration time depends on the system’s Hamiltonian, its initial state, and how it interacts with the heat bath. Its theoretical computation is very challenging, lying beyond the scope of the present work. Experimentally, it can be determined by perturbing the system appropriately using the above gates and measuring how long it takes to return to its equilibrium state.